# Cross-Modal Causal Intervention for Medical Report Generation

### v4

@@ IV-D3 Effectiveness of Causal Intervention
VDM. In Table VII, Baseline and R2Gen can boost the performance compared to Baseline and R2Gen, which demonstrates the validity of the VDM. However, the improvement of BLEU-4 between Baseline and Baseline on the IU-Xray dataset is more significant than that on the MIMIC-CXR dataset. This is because the VDM can discover essential visual information, but the report of the MIMIC-CXR is more complex and the model fails to generate accurate descriptions. The performance degradation of CIDEr can further illustrate it.
In Fig. 8, the encoder of the non-causal model exhibits limited attention to all potential abnormal regions. Instead, it excessively focuses on the base of the lung, possibly due to the dataset’s high prevalence of lung-related diseases. In contrast, the attention map from the VLCI encoder can truly focus on the dominant regions that may indicate abnormalities, including the entire lung lobe, carina, and pleura, rather than false correlations with biased visual concepts. This confirms the semantic sensitivity of the VDM, which captures dominant visual content by performing visual causal interventions.
LDM. Compared to the VDM, the LDM plays a more significant role in MRG because the sophisticated linguistic semantic patterns within reports are entangled and biased that require elaborate linguistic deconfounding. In Table VII, the performance drops without LDM, e.g., 0.119 0.110 for the BLEU-4 metric on the MIMIC-CXR dataset. This shows the importance of adjusting semantic relevance in word embedding space. Compared with the baseline, the performance improvement of Baseline on the MIMIC-CXR dataset demonstrates that the LDM can generate more accurate reports even with biased visual information. However, the CIDEr metric on the IU-Xray dataset shows the effectiveness of the combination of VDM and LDM, while ILVD obtains a lower score. This is due to the worse diversity on the IU-Xray dataset, where Baseline and R2Gen can get higher CIDEr but lower BLEU-4 with inadequate multi-modal feature correlation. In Fig. 8 (b), the attention map of the decoder in the non-causal model exhibits evident redundant responses, with high attention widely distributed in the upper part of the lung, especially the carina. The VLCI, in contrast, can capture dominant semantic information in a coarse-to-fine manner, refining it from the potential abnormal regions that receive extensive attention in the encoder to the bilateral thorax. The high attention on the carina may be attributed to the presence of a support device that could increase cardiac load, cause vascular occlusion or congestion, leading to changes in intrathoracic pressure and eventually resulting in pleural effusion. These findings indicate that the LDM can capture more discriminative semantic information from the linguistic modality through linguistic front-door interventions.
@@ V Conclusion
In this paper, we propose Visual-Linguistic Causal Intervention (VLCI) framework for MRG, to implicitly deconfound the visual-linguistic confounders by causal front-door intervention. To alleviate the problem of unpaired visual-linguistic data when pre-training, we combine the PLM and MIM for cross-modal pre-training. To implicitly mitigate cross-modal confounders and discover the true cross-modal causality, we propose visual-linguistic causal front-door intervention modules VDM and LDM. Experiments on IU-Xray and MIMIC-CXR datasets show that our VLCI can effectively mitigate visual-linguistic bias and outperforms the state-of-the-art methods. The lower computational cost and faster inference speed of VLCI promote its clinical application. In the future work, we will distill medical knowledge from large language models (LLMs) for more medical imaging modalities such as CT, MRI, and pathological slides. We believe our work could inspire more causal reasoning methods in MRG.

### v5

@@ IV-D3 Effectiveness of Causal Intervention
VDM. In Table VII, Baselinew◆ and R2Genw◆ can boost the performance compared to Baseline and R2Gen, which demonstrates the validity of the VDM. However, the improvement of BLEU-4 between Baselinew★◆ and Baselinew★ on the IU-Xray dataset is more significant than that on the MIMIC-CXR dataset. This is because the VDM can discover essential visual information, but the report of the MIMIC-CXR is more complex and the model fails to generate accurate descriptions. The performance degradation of CIDEr can further illustrate it.
In Fig. 11, the encoder of the non-causal model exhibits limited attention to all potential abnormal regions. Instead, it excessively focuses on the base of the lung, possibly due to the dataset’s high prevalence of lung-related diseases. In contrast, the attention map from the CMCRL encoder can truly focus on the dominant regions that may indicate abnormalities, including the entire lung lobe, carina, and pleura, rather than false correlations with biased visual concepts. This confirms the semantic sensitivity of the VDM, which captures dominant visual content by performing visual causal interventions.
LDM. Compared to the VDM, the LDM plays a more significant role in RRG because the sophisticated linguistic semantic patterns within reports are entangled and biased that require elaborate linguistic deconfounding. In Table VII, the performance drops without LDM, e.g., 11.9 11.0 for the BLEU-4 metric on the MIMIC-CXR dataset. This shows the importance of adjusting semantic relevance in word embedding space. Compared with the baseline, the performance improvement of Baselinew★∙ on the MIMIC-CXR dataset demonstrates that the LDM can generate more accurate reports even with biased visual information. However, the CIDEr metric on the IU-Xray dataset shows the effectiveness of the combination of VDM and LDM, while ILVD obtains a lower score. This is due to the worse diversity on the IU-Xray dataset, where Baselinew◆∙ and R2Genw◆∙ can get higher CIDEr but lower BLEU-4 with inadequate multi-modal feature correlation. In Fig. 11 (b), the attention map of the decoder in the non-causal model exhibits evident redundant responses, with high attention widely distributed in the upper part of the lung, especially the carina. The CMCRL, in contrast, can capture dominant semantic information in a coarse-to-fine manner, refining it from the potential abnormal regions that receive extensive attention in the encoder to the bilateral thorax. The high attention on the carina may be attributed to the presence of a support device that could increase cardiac load, cause vascular occlusion or congestion, leading to changes in intrathoracic pressure and eventually resulting in pleural effusion. These findings indicate that the LDM can capture more discriminative semantic information from the linguistic modality through linguistic front-door interventions.
@@ V Conclusion
In this paper, we propose the Cross-Modal Causal Representation Learning (CMCRL) framework for RRG, to deconfound visual-linguistic confounders by causal intervention. To alleviate the problem of unpaired visual-linguistic data when pre-training, we propose RadCARE for cross-modal pre-training, integrating postfix text generation and degradation-aware masked image restoration task. To mitigate cross-modal confounders and discover the true cross-modal causality, we propose visual-linguistic causal front-door intervention modules VDM and LDM, which are integrated into the visual-Linguistic Causal Intervention (VLCI) model. Experiments on IU-Xray and MIMIC-CXR datasets show that our CMCRL can effectively mitigate visual-linguistic bias and outperforms the state-of-the-art methods. The lower computational cost and faster inference speed of CMCRL promote its clinical application. In future work, we aim to further refine our method to better align with advanced Large Language Models (LLMs) and apply it to a broader range of medical imaging modalities, including CT, MRI, and digital pathology slides, among others. Additionally, to enhance the interpretability of causal reasoning, we will integrate the inferential capabilities of LLMs with their internal knowledge for causal inference. This may involve leveraging existing multimodal data to construct retrieval-augmented generation (RAG) tools or employing multi-agent systems to mine patient causal information from multidimensional medical records. We believe our work will inspire more causal reasoning methods within the realm of Radiology Report Generation (RRG) and extend its impact to other areas of medical imaging.